Let A=(2,-7) and B=(8,5). Find the equation of the line that contains the midpoint of AB and is perpendicular to AB. You may write your equation in either general or slope-intercept form.
I figured out the midpoint but I don't know where to go from there. The midpoint is (-3,-6). What next?
Thank you for your help!!!

Answer by gems(12)

the midpoint of x ( 2+8)/2=10/2=5
the midpoint of y ( -7+5)/2=-2/2=-1
so the midpoint of AB = ( 5, -1 ), NOT (-3,-6)
get the lope of AB
5--7/8-2
5+7/6
12/6
2
the slope is 2
get the negative reciprocal of 2,
-1/2
using point - slope form;
Y--1=-1/2(X-5)
Y+1=-1/2(X-5)
2(Y+1)=X-5
2Y+2=X-5
2Y-X+2+5=0
2Y-X+7=0
0R X-2Y-7=0
OR X-2Y=7